# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

### v2

@@ Appendix C Representation theory of $S_{N+1}$
The ($N+1$)-component vectors transform naturally under $\tilde{P}^{\alpha\beta}\in S_{N+1}$, where $\tilde{P}^{\alpha\beta}$ is a permutation matrix. However, this does not lead to an irreducible representation, since the ($N+1$)-component vector $(1,1,\ldots,1)$ forms the basis for a 1D invariant subspace. To construct a nontrivial irreducible representation, it is convenient to think of $S_{N+1}$ as the symmetry group of an $N$-simplex, which has $N+1$ vertices. The vertices are in one-to-one correspondence with the vectors $e_{a}^{\alpha}$ which connect the center and the vertices. The Latin index $a$ ranges from 1 to $N$, as the simplex is embedded in an $N$-dimensional space. The Greek index $\alpha$ ranges from 1 to $N+1$ because there are $N+1$ vertices. The vectors satisfy the relations Zia:1975ha  [^33] The normalization in Zia:1975ha is different from ours: $e_{a,\text{ours}}^{\alpha}=\sqrt{N}\;e_{a,\text{theirs}}^{\alpha}$.
@@
An $N$-component vector $v_{a}$ can be projected onto the vectors $e_{a}^{1},e_{a}^{2},\ldots,e_{a}^{N+1}$ (see figure 14), and the results can be viewed as an $(N+1)$-component vector which transforms under $\tilde{P}^{\alpha\beta}$. Then, we write the vector after the transformation as an $N$-component vector again. In short, the transformation rule for $N$-component vectors is given by
@@
In this way, the space of the $N$-component vectors furnishes an irreducible representation of $S_{N+1}$, which is isomorphic to the standard representation of $S_{N+1}$. [^34] The standard representation of $S_{N+1}$ consists of ($N+1$)-component vectors $u^{\alpha}$. They subject to the condition that the sum of the components vanishes $\sum_{\alpha=1}^{N+1}u^{\alpha}=0$. We do not distinguish between the two isomorphic vector spaces, and refer to the space of $N$-component vectors as the standard representation.
@@
Comparing the dimensions with those of (128), (129), (130) and (131), we conclude that (132), (133), (134) and (135) correspond to the Young diagrams $(N+1)$, $(N,1)$, $(N-1,2)$ and $(N-1,1^{2})$. [^35] The dimension of (134) follows from the symmetric tracelessness and the constraints $\sum_{a,b=1}^{N}T_{iab}\sum_{c,d=1}^{N}t^{(3)}_{abcd}\phi_{c}\phi_{d}=0\,,$ (137) where $i=1,2,\ldots,N$. From these examples, we see that the tensor product decomposition resembles that in the O($N$) model, except that here the symmetric traceless part further decomposes into two irreducible representations $V_{(N,1)}$ and $V_{(N-1,2)}$. Since $V_{(N,1)}$ is the standard representation, we use the term “symmetric traceless representation” to refer to $V_{(N-1,2)}$. Moreover, we refer to $V_{(N+1)}$ and $V_{(N-1,1^{2})}$ as the “trivial representation” and the “antisymmetric representation”. [^36] The antisymmetric representation $V_{(N-1,1^{2})}$ is the exterior power $\Lambda^{2}V_{(N,1)}$ of the standard representation $V_{(N,1)}$.
@@ Appendix D Tensor structures
The tensor product decomposition of $(N,1)^{\otimes 4}$ contains the trivial representation four times, so there are four linearly independent rank-4 invariant tensors wallace1978spin . Using $e_{a}^{\alpha}$ as defined in (126), we identify four rank-4 $S_{N+1}$-invariant tensors: $\delta_{ab}\delta_{cd}$, $\delta_{ac}\delta_{bd}$, $\delta_{ad}\delta_{bc}$ and $\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}$. We then use these tensors to define a basis for the four-point function $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\phi_{c}(x_{3})\phi_{d}(x_{4})\rangle$:
@@
For $i=1,2,3,4$, the tensor structures $t_{abcd}^{(i)}$ are associated with the intermediate operators in the trivial, standard, symmetric tracless and antisymmetric representations. The four tensors can also be used as projectors, which satisfy
@@
The normalization of $t_{abcd}^{(i)}$ is chosen as follows. The free correlator $\langle\phi_{a,\text{f}}(x_{1})\phi_{b,\text{f}}(x_{2})\phi_{c,\text{f}}(x_{3})$ $\phi_{d,\text{f}}(x_{4})\rangle$ is computed using Wick’s theorem, yielding four parts that correspond to the four tensor structures. We require that the conformal block decomposition of each part is precisely (34), with the same normalization of OPE coefficients and conformal blocks. We use these projectors to construct bilinear operators in the irreducible representations, as in (132), (133), (134) and (135).
@@
$+\frac{(N-2)(N+1)}{2N^{2}}g^{(3)}(1-\bar{z},1-z)+\frac{1-N}{2N}g^{(4)}(1-\bar{z},1-z)\Bigg{)}\,,$ (144)
@@
$-\frac{N-2}{2N(N+1)(N-1)^{2}}g^{(3)}(1-\bar{z},1-z)+\frac{g^{(4)}(1-\bar{z},1-z)}{2(N-1)(N+1)^{2}}\Bigg{)}\,,$ (145)
@@
$+\frac{N^{2}-N+2}{2N(N-1)}g^{(3)}(1-\bar{z},1-z)+\frac{g^{(4)}(1-\bar{z},1-z)}{2}\Bigg{)}\,,$ (146)
@@
$+\frac{(N-2)(N+1)}{2N(N-1)}g^{(3)}(1-\bar{z},1-z)+\frac{g^{(4)}(1-\bar{z},1-z)}{2}\Bigg{)}\,.$ (147)
The analytic bootstrap in subsection 3.2 is based on these crossing equations.
@@ Appendix E Ratios of OPE coefficients
Let us first focus on the $N=1$ case. We only need the free theory data here, and we can consider the ratios with $n=1$ and $n>1$ together. The four-point function of identical scalars reads
@@
where the factors $(2n)!$ and $(2n)(2n)!$ come from Wick contraction. The kinematic factors are given by
@@
=\;= $\Bigg{[}\left(1+\frac{\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)}{N}\right)t_{abcd}^{(1)}$
$\quad+\frac{\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)}{(N-1)(N+1)^{2}}t_{abcd}^{(2)}$
$\quad+\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)t_{abcd}^{(3)}$
$\quad+\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1-\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)t_{abcd}^{(4)}\Bigg{]}\mathcal{K}(\Delta_{i},x_{i})$ (153)
@@
Appendix F An OPE coefficient in generalized $\phi^{2n+1}$ theory ($n>1$)
We are interested in the OPE coefficient $\lambda_{\phi\phi\phi^{2n-1}}$ in $\phi\times\phi\sim\phi^{2n-1}$. This coefficient is useful in the analytic bootstrap, as in (4.2). We consider the matching condition
@@ Appendix F An OPE coefficient in generalized $\phi^{2n+1}$ theory ($n>1$)
$\lambda_{\phi\phi\phi^{2n-1}}=\frac{(2n)!\,\alpha}{2^{2k}(1-2k)_{k}\left(k\frac{3-2n}{2n-1}\right)_{k}}+O(\alpha^{2})\,.$ (159)
According to the convention (37) in the analytic bootstrap, we obtain
@@
Appendix G The $N\rightarrow 0$ limit, logarithmic CFTs and logarithmic recombination
As mentioned in footnote 8, the $N\rightarrow 0$ limit of the Potts model corresponds to a logarithmic CFT. A key feature is that the tensor structure (140) diverges in the $N\rightarrow 0$ limit. This singularity can be resolved by mixing bilinear operators. We will see below that this mixing leads to logarithmic structures and nonvanishing off-diagonal two-point functions. This means that the primaries are not independent and form logarithmic multiplets, which can be viewed as a logarithmic version of the multiplet recombination. The discussion in this appendix applies to the canonical case $k=1$ and the higher-derivative cases with $k>1$.
In the above, the tensor structure (140) is associated with the two-point functions of bilinear operators $\mathcal{J_{\text{ST}}}$ in the symmetric traceless (ST) representation. To cancel the $1/N$ pole, $\mathcal{J_{\text{ST}}}$ should mix with $\mathcal{J_{\text{trivial}}}$ and they form new operators.373737 Although the tensor structure (140) is also singular in the $N\rightarrow-1$ limit, the two-point function of $\phi^{2}_{\text{ST}}$ vanishes due to the normalization (see (162)). At order $\epsilon^{0}$, the OPE coefficient vanishes as well (see (169)), but it can be made finite by a singular change in the normalization $\bar{\phi}^{2}=\frac{1}{N+1}\phi^{2}_{\text{ST}}$. For instance, the two-point functions of the spin-0 operators with $m=0$ take the form
@@ Appendix G The $N\rightarrow 0$ limit, logarithmic CFTs and logarithmic recombination
Here the normalization of the two-point functions is the same as that in the free theory, i.e., the two-point functions are given by Wick’s theorem and $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\rangle=|x_{1}-x_{2}|^{-2\smash{\Delta_{\phi}}}\delta_{ab}$. The explicit expressions for $\phi^{2}_{ab,\text{ST}}$ and $\phi^{2}_{\text{trivial}}$ can be found in (132) and (134). A linear combination that is regular in the $N\rightarrow 0$ limit reads
@@
$\hskip 30.00005pt-2(N+1)(N-1)\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}+\frac{8\epsilon}{7-3N}\log x_{12}\,\delta_{ab}\delta_{cd}+O(\epsilon^{3/2})\Bigg{)}\Bigg{]}$
=\;= $\frac{1}{x_{12}^{2\Delta_{0}}}\left(-2\delta_{ab}\delta_{cd}+2\left(\delta_{ac}\delta_{bd}+\delta_{ad}\delta_{bc}\right)+2\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}+\frac{8\epsilon}{7}\log x_{12}\,\delta_{ab}\delta_{cd}+O(\epsilon^{3/2})\right)\,.$ (164)
@@
=\;= $\frac{1}{x_{12}^{2\Delta_{0}}}\begin{pmatrix}\bar{t}_{abcd}+\frac{8\epsilon}{7}\log x_{12}\,t^{(1)}_{abcd}&\;-2t^{(1)}_{abcd}\\ -2t^{(1)}_{abcd}&0\end{pmatrix}+O(\epsilon^{3/2})\,,$ (166)
@@
We interpret the matrix elements as the two-point functions of the primaries in a rank-2 logarithmic multiplet. Although the norm of $\phi^{2}_{\text{trivial}}$ vanishes in the $N\rightarrow 0$ limit, $\phi^{2}_{\text{trivial}}$ is not a null state and have nonzero correlation functions with other operators.
Let us verify that the three-point function $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\bar{\phi}^{2}_{cd}(x_{3})\rangle$ is regular as well. We consider the three-point functions of two $\phi$’s and a bilinear operator
@@
Here $h_{ijk}=\Delta_{i}+\Delta_{j}-\Delta_{k}$, and $\Delta_{i}$ is the scaling dimension of the operator at position $x_{i}$. The leading term in the OPE coefficient for $\phi^{2}_{\text{trivial}}$ does not depend on $N$, so it remains finite in the $N\rightarrow 0$ limit, as mentioned above. However, the tensor structure $t^{(3)}_{abcd}$ in the second line diverges as $N$ approaches zero. This singular behaviour is also resolved by the linear combination (163), yielding the three-point function
@@
According to the results for the correlation functions, the two independent primaries $\phi^{2}_{\text{trivial}}$ and $\phi^{2}_{\text{ST}}$ mix and form the logarithmic primary $\bar{\phi}^{2}_{\text{ab}}$ in the limit $N\rightarrow 0$. The two primary states associated with $\bar{\phi}^{2}_{\text{ab}}$ and $\phi^{2}_{\text{trivial}}$ are not orthogonal to each other, giving rise to nonvanishing off-diagonal two-point functions. In other words, $\{\phi^{2}_{\text{trivial}}\}$ and $\{\bar{\phi}^{2}_{\text{ab}}\}$ recombine into a logarithmic multiplet. As the two multiplets are related by the derivative with respective to the scaling dimension $\Delta$, the logarithmic recombination is different from the multiplet recombination based on the derivative with respective to $x^{\mu}$ in the main text. The logarithmic version of the multiplet recombination may deserve further investigations.
@@
$+\left(\frac{1}{3+\frac{(-1)^{k}k!(3k)_{k}}{2^{4k-5}(1/2)_{k}^{2}}}+\frac{(2k)!(2k-m)_{k+m}}{\left(8k!+\frac{3\times 2^{4k-2}(-1)^{k}(1/2)_{k}^{2}}{(3k)_{k}}\right)(1-3k-\ell-m)_{2k}(-m)_{m}}\right)\epsilon+O(\epsilon^{3/2})\,.$ (171)
@@
$\frac{d-\Delta_{\phi^{2}_{\text{ST}}}|_{N\rightarrow 0}}{d-\Delta_{\phi^{2}_{\text{trivial}}}|_{N\rightarrow 0}}=1+O(\epsilon^{3/2})\,.$ (172)
@@
There is a similar story in the $N\rightarrow 0$ limit of the generalized O($N$)-symmetric $\phi^{4}$ theory. For general $N$, the scalar two-point functions of singlet (S) and symmetric traceless (ST) bilinear operators are
@@
As in (163), the coefficient in front of $\phi^{2}_{\text{S}}$ is not unique. Furthermore, we expect that the degeneracies extend to the spinning operators. The scaling dimensions of the singlet and symmetric traceless bilinear operators indeed become identical in the $N\rightarrow 0$ limit:
$\Delta_{\mathcal{J}_{\text{S}}}\big{|}_{N\rightarrow 0}=\Delta_{\mathcal{J}_{\text{ST}}}\big{|}_{N\rightarrow 0}$
=\;= $d-2k+2m+\ell-\frac{(k-1)!(k+1)_{k}^{2}}{128(k)_{k-m}}\left(\frac{1}{(-k)_{k}(2k-m)_{k+m}}+\frac{3}{(-m)_{m}(m+\ell)_{2k}}\right)\epsilon^{2}+O(\epsilon^{3})\,.$ (176)
@@
Appendix H Generalized self-avoiding walks, loop-erased random walks and their $\phi^{2n}$ generalizations
As prominent models of random walks, we consider the self-avoiding walks and loop-erased random walks that play important roles in physics. Interestingly, they can be associated with the O($N$) symmetric $\phi^{4}$ theory. The self-avoiding walks correspond to the $N\rightarrow 0$ limit deGennes:1972zz  [^38] See Liu:2012ca for some Monte Carlo results for the loop model at various $N$ in three dimensions, including the case of $N=0$. and the loop-erased random walks are related to the $N\rightarrow-2$ limit Nienhuis:1982fx ; duplantier1992loop ; Wiese:2018dow ; Wiese:2019xmu ; helmuth2020loop . In this appendix, we discuss these limits and their $\phi^{2n}$ generalizations using the results from our previous paper Guo:2023qtt .
In Guo:2023qtt , we studied the generalized $\phi^{2n}$ theory with O($N$) global symmetry. The anomalous dimensions of $\phi$ at $n=2$ reads

### v3

@@ Appendix C Representation theory of $S_{N+1}$
The ($N+1$)-component vectors transform naturally under $\tilde{P}^{\alpha\beta}\in S_{N+1}$, where $\tilde{P}^{\alpha\beta}$ is a permutation matrix. However, this does not lead to an irreducible representation, since the ($N+1$)-component vector $(1,1,\ldots,1)$ forms the basis for a 1D invariant subspace. To construct a nontrivial irreducible representation, it is convenient to think of $S_{N+1}$ as the symmetry group of an $N$-simplex, which has $N+1$ vertices. The vertices are in one-to-one correspondence with the vectors $e_{a}^{\alpha}$, which connect the center and vertices. The latin index $a$ ranges from 1 to $N$ as the simplex is embedded in an $N$-dimensional space. The greek index $\alpha$ ranges from 1 to $N+1$ because there are $N+1$ vertices. The vectors satisfy the relations Zia:1975ha  [^41] The normalization in Zia:1975ha is different from ours: $e_{a,\text{ours}}^{\alpha}=\sqrt{N}\;e_{a,\text{theirs}}^{\alpha}$.
@@
An $N$-component vector $v_{a}$ can be projected onto the vectors $e_{a}^{1},e_{a}^{2},\ldots,e_{a}^{N+1}$ (see figure 14), and the results can be viewed as an $(N+1)$-component vector that transforms under $\tilde{P}^{\alpha\beta}$. Then, we write the vector after the transformation as an $N$-component vector again. In short, the transformation rule for $N$-component vectors is given by
@@
In this way, the space of the $N$-component vectors furnishes an irreducible representation of $S_{N+1}$, which is isomorphic to the standard representation of $S_{N+1}$. [^42] The standard representation of $S_{N+1}$ consists of ($N+1$)-component vectors $u^{\alpha}$. They are subject to the condition that the sum of the components vanishes $\sum_{\alpha=1}^{N+1}u^{\alpha}=0$. We do not distinguish between the two isomorphic vector spaces and refer to the space of $N$-component vectors as the standard representation.
@@
Comparing the dimensions with those of (130), (131), (132), and (133), we conclude that (134), (135), (136), and (137) correspond to the Young diagrams $(N+1)$, $(N,1)$, $(N-1,2)$, and $(N-1,1^{2})$. [^43] The dimension of (136) follows from the symmetric tracelessness and the constraints $\sum_{a,b=1}^{N}T_{iab}\sum_{c,d=1}^{N}t^{(3)}_{abcd}\phi_{c}\phi_{d}=0\,,$ (139) where $i=1,2,\ldots,N$. From these examples, we see that the tensor product decomposition resembles that in the O($N$) model, except that here the symmetric traceless part further decomposes into two irreducible representations $V_{(N,1)}$ and $V_{(N-1,2)}$. Since $V_{(N,1)}$ is the standard representation, we use the term “symmetric traceless representation” to refer to $V_{(N-1,2)}$. Moreover, we refer to $V_{(N+1)}$ and $V_{(N-1,1^{2})}$ as the “trivial representation” and the “antisymmetric representation.” [^44] The antisymmetric representation $V_{(N-1,1^{2})}$ is the exterior power $\Lambda^{2}V_{(N,1)}$ of the standard representation $V_{(N,1)}$.
@@ Appendix D Tensor structures
The tensor product decomposition of $(N,1)^{\otimes 4}$ contains the trivial representation four times, so there are four linearly independent rank-4 invariant tensors wallace1978spin . Using $e_{a}^{\alpha}$ as defined in (128), we identify four rank-4 $S_{N+1}$-invariant tensors: $\delta_{ab}\delta_{cd}$, $\delta_{ac}\delta_{bd}$, $\delta_{ad}\delta_{bc}$, and $\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}$. We then use these tensors to define a basis for the four-point function $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\phi_{c}(x_{3})\phi_{d}(x_{4})\rangle$:
@@
For $i=1,2,3,4$, the tensor structures $t_{abcd}^{(i)}$ are associated with the intermediate operators in the trivial, standard, symmetric tracless, and antisymmetric representations. The four tensors can also be used as projectors, which satisfy
@@
The normalization of $t_{abcd}^{(i)}$ is chosen as follows. The free correlator $\langle\phi_{a,\text{f}}(x_{1})\phi_{b,\text{f}}(x_{2})\phi_{c,\text{f}}(x_{3})$ $\phi_{d,\text{f}}(x_{4})\rangle$ is computed using Wick contractions, yielding four parts that correspond to the four tensor structures. We require that the conformal block decomposition of each part is precisely (32), with the same normalization of OPE coefficients and conformal blocks. We use these projectors to construct bilinear operators in the irreducible representations, as in (134), (135), (136), and (137).
@@
$\hskip 60.00009pt+\frac{(N-2)(N+1)}{2N^{2}}g^{(3)}(1-\bar{z},1-z)+\frac{1-N}{2N}g^{(4)}(1-\bar{z},1-z)\Bigg{)}\,,$ (146)
@@
$\hskip 60.00009pt-\frac{N-2}{2N(N+1)(N-1)^{2}}g^{(3)}(1-\bar{z},1-z)+\frac{g^{(4)}(1-\bar{z},1-z)}{2(N-1)(N+1)^{2}}\Bigg{)}\,,$ (147)
@@
$\hskip 100.00015pt+\frac{N^{2}-N+2}{2N(N-1)}g^{(3)}(1-\bar{z},1-z)+\frac{g^{(4)}(1-\bar{z},1-z)}{2}\Bigg{)}\,,$ (148)
@@
$\hskip 90.00014pt+\frac{(N-2)(N+1)}{2N(N-1)}g^{(3)}(1-\bar{z},1-z)+\frac{g^{(4)}(1-\bar{z},1-z)}{2}\Bigg{)}\,.$ (149)
The analytic bootstrap in section 3.2 is based on these crossing equations.
@@ Appendix E Ratios of OPE coefficients
Let us first focus on the $N=1$ case. We only need the free-theory data here, and we can consider the ratios with $n=1$ and $n>1$ together. The four-point function of identical scalars reads
@@
where the factors $(2n)!$ and $(2n)(2n)!$ come from Wick contractions. The kinematic factors are given by
@@
=\;= $\Bigg{[}\left(1+\frac{\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)}{N}\right)t_{abcd}^{(1)}+\frac{\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)}{(N-1)(N+1)^{2}}t_{abcd}^{(2)}$
$+\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)t_{abcd}^{(3)}$
$+\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1-\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)t_{abcd}^{(4)}\Bigg{]}\mathcal{K}(\Delta_{i},x_{i})$ (155)
@@
Appendix F An OPE coefficient in generalized $\phi^{2n+1}$ theory
We are interested in the OPE coefficient $\lambda_{\phi\phi\phi^{2n-1}}$. This coefficient is useful in the analytic bootstrap, as in (4.2). We consider the matching condition
@@ Appendix F An OPE coefficient in generalized $\phi^{2n+1}$ theory
$\lambda_{\phi\phi\phi^{2n-1}}=\frac{(-1)^{k}(2n)!\,\alpha}{2^{2k}(k)_{k}\left(-k\frac{2n-3}{2n-1}\right)_{k}}+\ldots\,,$ (161)
which corresponds to (22) at $n=1$. According to the convention (35) in the analytic bootstrap, we obtain
@@
Appendix G The $N\rightarrow 0$ limit, logarithmic CFTs, and logarithmic recombination
As mentioned in footnote 10, the $N\rightarrow 0$ limit of the Potts model corresponds to a logarithmic CFT. A key feature is that the tensor structure (142) diverges in the $N\rightarrow 0$ limit. This singularity can be resolved by mixing bilinear operators. We will see below that this mixing leads to logarithmic structures and nonvanishing off-diagonal two-point functions. This means that the primaries are not independent and form logarithmic multiplets, which can be viewed as a logarithmic version of the multiplet recombination. The discussion in this appendix applies to the canonical case $k=1$ and the higher-derivative cases with $k>1$.
In the above, the tensor structure (142) is associated with the two-point functions of bilinear operators $\mathcal{J_{\text{ST}}}$ in the symmetric traceless (ST) representation. To cancel the $1/N$ pole, $\mathcal{J_{\text{ST}}}$ should mix with $\mathcal{J_{\text{trivial}}}$ and they form new operators.454545 Although the tensor structure (142) is also singular in the $N\rightarrow-1$ limit, the two-point function of $\phi^{2}_{\text{ST}}$ vanishes due to the normalization [see (164)]. At order $\epsilon^{0}$, the OPE coefficient vanishes as well [see (171)], but it can be made finite by a singular change in the normalization $\bar{\phi}^{2}=\frac{1}{N+1}\phi^{2}_{\text{ST}}$. For instance, the two-point functions of the spin-0 operators with $m=0$ take the form
@@ Appendix G The $N\rightarrow 0$ limit, logarithmic CFTs, and logarithmic recombination
Here the normalization of the two-point functions is the same as that in the free theory, i.e., the two-point functions are given by Wick contractions and $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\rangle=|x_{1}-x_{2}|^{-2\smash{\Delta_{\phi}}}\delta_{ab}$. The explicit expressions for $\phi^{2}_{ab,\text{ST}}$ and $\phi^{2}_{\text{trivial}}$ can be found in (134) and (136). A linear combination that is regular in the $N\rightarrow 0$ limit reads
@@
$\hskip 30.00005pt-2(N+1)(N-1)\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}+\frac{8\epsilon}{7-3N}\log x_{12}\,\delta_{ab}\delta_{cd}+\ldots\Bigg{)}\Bigg{]}$
=\;= $\frac{1}{x_{12}^{2\Delta_{0}}}\left(-2\delta_{ab}\delta_{cd}+2\left(\delta_{ac}\delta_{bd}+\delta_{ad}\delta_{bc}\right)+2\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}+\frac{8\epsilon}{7}\log x_{12}\,\delta_{ab}\delta_{cd}+\ldots\right)\,.$ (166)
@@
=\;= $\frac{1}{x_{12}^{2\Delta_{0}}}\begin{pmatrix}\bar{t}_{abcd}+\frac{8\epsilon}{7}\log x_{12}\,t^{(1)}_{abcd}&\;-2t^{(1)}_{abcd}\\ -2t^{(1)}_{abcd}&0\end{pmatrix}+\ldots\,,$ (168)
@@
We interpret the matrix elements as the two-point functions of the primaries in a rank-2 logarithmic multiplet. Although the norm of $\phi^{2}_{\text{trivial}}$ vanishes in the $N\rightarrow 0$ limit, $\phi^{2}_{\text{trivial}}$ is not a null state and has nonzero correlation functions with other operators.
Let us verify that the three-point function $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\bar{\phi}^{2}_{cd}(x_{3})\rangle$ is regular as well. We consider the three-point functions of two $\phi$’s and a bilinear operator,
@@
Here $h_{ijk}=\Delta_{i}+\Delta_{j}-\Delta_{k}$, and $\Delta_{i}$ is the scaling dimension of the operator at position $x_{i}$. The leading term in the OPE coefficient for $\phi^{2}_{\text{trivial}}$ does not depend on $N$, so it remains finite in the $N\rightarrow 0$ limit, as mentioned above. However, the tensor structure $t^{(3)}_{abcd}$ in the second line diverges as $N$ approaches zero. This singular behavior is also resolved by the linear combination (165), yielding the three-point function
@@
According to the results for the correlation functions, the two independent primaries $\phi^{2}_{\text{trivial}}$ and $\phi^{2}_{\text{ST}}$ mix and form the logarithmic primary $\bar{\phi}^{2}_{\text{ab}}$ in the limit $N\rightarrow 0$. The two primary states associated with $\bar{\phi}^{2}_{\text{ab}}$ and $\phi^{2}_{\text{trivial}}$ are not orthogonal to each other, giving rise to nonvanishing off-diagonal two-point functions. In other words, $\{\phi^{2}_{\text{trivial}}\}$ and $\{\bar{\phi}^{2}_{\text{ab}}\}$ recombine into a logarithmic multiplet. As the two multiplets are related by the derivative with respect to the scaling dimension $\Delta$, the logarithmic recombination is different from the multiplet recombination based on the derivative with respect to $x^{\mu}$ in the main text. The logarithmic version of the multiplet recombination may deserve further investigations.
@@
$\hskip 50.00008pt+\left(\frac{1}{3+\frac{(-1)^{k}k!(3k)_{k}}{2^{4k-5}(1/2)_{k}^{2}}}+\frac{(-1)^{m}(2k)!(2k-m)_{k+m}}{\left(8k!+\frac{3\times 2^{4k-2}(-1)^{k}(1/2)_{k}^{2}}{(3k)_{k}}\right)m!(k+\ell+m)_{2k}}\right)\epsilon+\ldots$ (173)
@@
$\frac{d-\Delta_{\phi^{2}_{\text{ST}}}|_{N\rightarrow 0}}{d-\Delta_{\phi^{2}_{\text{trivial}}}|_{N\rightarrow 0}}=1+0\epsilon+\ldots\,.$ (174)
@@
There is a similar story in the $N\rightarrow 0$ limit of the generalized O($N$)-symmetric $\phi^{4}$ theory. For general $N$, the scalar two-point functions of singlet (S) and ST bilinear operators are
@@
As in (165), the coefficient in front of $\phi^{2}_{\text{S}}$ is not unique. Furthermore, we expect that the degeneracies extend to the spinning operators. The scaling dimensions of the S and ST bilinear operators indeed become identical in the $N\rightarrow 0$ limit:
$\Delta_{\mathcal{J}_{\text{S}}}\big{|}_{N\rightarrow 0}=\Delta_{\mathcal{J}_{\text{ST}}}\big{|}_{N\rightarrow 0}=\;$ $d-2k+2m+\ell$
$-\frac{(k-1)!(k+1)_{k}^{2}}{128(k)_{k-m}}\left(\frac{(-1)^{k}}{k!(2k-m)_{k+m}}+\frac{3(-1)^{m}}{m!(m+\ell)_{2k}}\right)\epsilon^{2}+O(\epsilon^{3})\,.$ (178)
@@
Appendix H Generalized self-avoiding walks, loop-erased random walks, and their $\phi^{2n}$ generalizations
As prominent models of random walks, we consider the self-avoiding walks and loop-erased random walks that play important roles in physics. Interestingly, they can be associated with the O($N$)-symmetric $\phi^{4}$ theory. The self-avoiding walks correspond to the $N\rightarrow 0$ limit deGennes:1972zz  [^46] See Liu:2012ca for some Monte Carlo results for the loop model at various $N$ in three dimensions, including the case of $N=0$. and the loop-erased random walks are related to the $N\rightarrow-2$ limit Nienhuis:1982fx ; duplantier1992loop ; Wiese:2018dow ; Wiese:2019xmu ; helmuth2020loop . In this appendix, we discuss these limits and their $\phi^{2n}$ generalizations using the results from our previous paper Guo:2023qtt .
In Guo:2023qtt , we studied the generalized $\phi^{2n}$ theory with O($N$) global symmetry. The anomalous dimension of $\phi$ at $n=2$ reads